# BarrierBench : Evaluating Large Language Models for Safety Verification in Dynamical Systems

### v1

@@ Abstract
Safety verification of dynamical systems via barrier certificates is central to ensuring correctness in autonomous and safety-critical applications. Synthesizing such certificates requires discovering mathematical functions that characterize inductive state invariants which provably separate safe and unsafe regions. However, existing approaches struggle with computational scalability, depend on careful template design, and often rely on exhaustive or incremental searches over function spaces. Moreover, these approaches demand significant manual ingenuity and mathematical sophistication in constructing the search infrastructure: selecting template structures, choosing solvers, tuning hyperparameters for data-driven methods, and designing effective sampling procedures. A successful synthesis of barrier certificates therefore requires both a deep understanding of system and control theory and practical experience with these techniques. This expertise has traditionally been passed among practitioners through natural language rather than being formalized mathematically.
This observation raises a natural question: can the linguistic and analogical reasoning that experts use informally be captured and operationalized by language models? Motivated by this idea, we present a Large Language Model (LLM) agentic framework for barrier certificate synthesis that leverages natural language reasoning to propose, refine, and validate candidate certificates. The framework integrates LLM-driven template discovery with SMT-based verification and supports barrier-controller co-synthesis for controlled systems, ensuring mathematical compatibility between safety certificates and feedback control laws.
To evaluate this capability, we introduce BarrierBench, a benchmark of 100 dynamical systems spanning linear, nonlinear, discrete-time, and continuous-time settings, including 68 controlled systems that require barrier–controller co-synthesis. Our experiments assess not only the effectiveness of LLM-guided barrier synthesis but also the utility of retrieval-augmented generation (RAG) and agentic coordination strategies in improving its reliability and performance. Across these tasks, the framework achieves more than 90% success in generating valid certificates and demonstrates structural diversity ranging from simple quadratics to high-order coupled polynomials. By releasing BarrierBench and the accompanying toolchain, we aim to establish a community testbed for advancing the integration of language-based reasoning with formal verification in dynamical systems. The benchmark is publicly available at https://hycodev.com/dataset/barrierbench.
@@ 1 Introduction
This century has witnessed an increasingly sophisticated integration of digital software with physical infrastructure, interconnected through wired and wireless networks. These cyber-physical systems (CPS) operate so seamlessly and intuitively that it is often easy to overlook the complexity beneath their surface: their success depends on intricate control logic whose correctness requires a deep understanding of both the underlying physics and the discrete switching logic that governs their operation. As a result, while technological practice continues to deliver critical infrastructures—such as autonomous vehicles, implantable medical devices, and smart grids—the task of control engineers to ensure their safety remains a Sisyphean endeavor amid ever-growing system complexity. In this paper, we address this enduring challenge by exploring how large language models can assist in the synthesis and verification of safety guarantees for complex CPS.
Safety Verification.
@@
Safety verification of dynamical systems is a canonical challenge in the design of safety-critical CPS. Since safety verification is computationally undecidable even for CPS with simple piecewise-constant dynamics (asarin1998achilles), achieving provable safety through fully automatic procedures remains out of reach. Consequently, a range of techniques has been developed to construct safety proofs, including symbolic state-space exploration, statistical guarantees, and abstraction-based methods. A particularly effective and scalable alternative is the synthesis of barrier certificates (prajna2004safety; ames2019control; jagtap2020formal), which serve as functional analogs of inductive state invariants and characterize the separation between safe and unsafe regions of the state space. With origins in Lyapunov’s classical theory of stability (lyapunov1966stability; khalil2002nonlinear), barrier certificates extend its principle from proving convergence to demonstrating invariance. Their formulation as a search for a single scalar function subject to inequality constraints makes them amenable to optimization and satisfiability-based methods, contributing to their popularity in both theory and practice. We focus on the barrier-certificate-based approach to proof discovery as a promising foundation for integrating formal reasoning with language-model-guided synthesis.
Barrier Certificates.
@@
Formally, a barrier certificate is a scalar function whose level sets define an inductive invariant that separates unsafe regions from all trajectories starting from admissible initial conditions. By providing formal guarantees over infinite time horizons, barrier certificates have become central to ensuring correctness in both linear and nonlinear dynamical systems. Traditional approaches to barrier synthesis are predominantly optimization-based. Sum-of-squares and semidefinite programming methods formulate the search as a convex optimization problem over polynomial templates (prajna2007framework; kordabad2025sum), yielding provable correctness but suffering from exponential complexity growth with system dimension and being restricted to low-degree polynomial classes. To improve generality, counterexample-guided inductive synthesis frameworks (see the work by (dawson2023safe) for a comprehensive review) extend these techniques using SMT solvers to iteratively refine candidate barriers (ravanbakhsh2015counter), whereas neural and data-driven variants employ deep architectures to approximate more expressive certificate functions (abate2024safe; neustroev2025neural; abate2021fossil). Although such methods broaden the space of representable barriers, they introduce new challenges—instability in training, reliance on finite samples, reduced interpretability, and a persistent need for careful template or architecture design. Consequently, successful synthesis continues to depend heavily on expert intuition: selecting suitable function classes, solvers, and hyperparameters. These choices often encode informal reasoning that remains difficult to capture or automate within existing algorithmic frameworks. This paper investigates whether large language models (LLMs) can capture and operationalize the linguistic and analogical reasoning that practitioners employ when constructing barrier certificates.
Language-Model-Guided Synthesis.
@@
We present an LLM-guided synthesis framework for barrier certificates that leverages natural-language reasoning to propose, refine, and validate candidate certificates. The framework integrates LLM-driven template discovery with SMT-based formal verification and extends naturally to barrier–controller co-synthesis for controlled systems. By incorporating verification feedback into subsequent LLM prompts, the framework enables dynamic exploration of barrier certificate templates and coefficients, often discovering novel mathematical forms that traditional fixed-template approaches (prajna2007framework; kong2013exponential; ames2019control; abate2021fossil; kordabad2025sum) might overlook. LLMs guide the synthesis process through reasoning about system properties, safety requirements, and mathematical structures, systematically analyzing local safety properties before integrating them into global certificates. This integration provides both the expressiveness and adaptability needed for complex systems while preserving the mathematical rigor essential for safety-critical applications.
Figure 1: Architecture of the agentic framework integrating Retrieval, Synthesis, and Verifier agents for LLM-guided barrier certificate synthesis.
@@
To operationalize this process, we design an agentic architecture that emulates the workflow of expert control designers. Human designers typically draw on prior experience with similar dynamical systems to select suitable barrier templates, solver strategies, and proof heuristics. Analogously, our framework instantiates three collaborating agents: a Barrier Retrieval Agent that searches a database of previously solved systems to identify analogous examples; a Barrier Synthesis Agent that proposes candidate barrier certificates inspired by retrieved analogs; and a Barrier Verifier Agent that evaluates these candidates in two stages—a lightweight, data-driven check followed by formal verification via an SMT solver. Feedback from the Verifier Agent is fed back to the Synthesis Agent, forming a closed-loop interaction among agents bounded by a global timeout. This design enables iterative refinement of candidate certificates and efficient use of external formal tools. Our study investigates the following research questions:
@@ 3. RQ3:
BarrierBench.
To evaluate these questions, we introduce BarrierBench, a benchmark of 100 dynamical systems spanning linear, nonlinear, discrete-time, and continuous-time settings, including 68 controlled systems that require barrier–controller co-synthesis. Across these tasks, the LLM framework achieves more than 90% success in generating valid certificates and exhibits structural diversity ranging from simple quadratics to high-order coupled polynomials. By releasing BarrierBench and the accompanying toolchain, we aim to establish a community testbed for advancing the integration of language-based reasoning with formal verification in dynamical systems.
@@ 2 Barrier Certificates
Theorem 1 (Safety via Barrier Certificates (prajna2004safety)).
@@
The above definitions and theorem are stated for systems without control inputs. If the system has control input $u\in U$, the objective is to design a state feedback that gives the input $u$ as a function of state $x$, which makes the system safe. For this, the conditions of barrier will be quantified existentially over the input.
@@
 $\text{example}\leftarrow\textsc{RetrievalAgent}(f,X_{I},X_{U},\mathcal{D})$; $(B_{k},C_{k})\leftarrow\textsc{SynthesisAgent}(\text{example},\textsf{feedback}_{k})$
@@ 3 Agentic Framework for Barrier Synthesis
Our synthesis framework employs a multi-agent architecture in which specialized LLM-powered agents collaborate to discover barrier certificates. The framework comprises three core agents operating within an iterative pipeline: (1) a Barrier Retrieval Agent that identifies relevant historical solutions; (2) a Barrier Synthesis Agent that generates candidate barrier certificates; and (3) a Barrier Verifier Agent that verifies candidates and provides feedback for progressive refinement. This iterative process overcomes limitations of conventional approaches by enabling intelligent template selection and adaptive exploration of diverse mathematical structures through verification-guided synthesis.
@@
The Barrier Verifier Agent conducts a two-stage evaluation of candidate barriers. In the first stage, it performs a sample-based validation across $N$ points drawn from $X_{I}$, $X_{U}$, and $X$, where $N$ is selected to balance computational efficiency with adequate state space coverage; we use $N=5000$ in our experiments to achieve this trade-off, discarding invalid barriers to avoid unnecessary SMT solver calls. In the second stage, it performs formal verification using SMT solvers such as Z3 (demoura2008z3), Yices (dutertre2006yices), and cvc5 (barbosa2022cvc5), checking the unsatisfiability of negated barrier conditions. The agent selects solvers adaptively based on problem structure, applies symbolic conversions or Taylor approximations when necessary, and handles verification failures through solver switching or adaptive timeout adjustments.
@@ 4.1 Experimental Setup
The framework operates with carefully chosen hyperparameters validated across various systems. We set maximum main iterations $K=5$ and maximum refinements per iteration $R=4$. The LLM utilizes Claude Sonnet 4 and ChatGPT-4o with engineered prompts designed to synthesize mathematical expressions while maintaining formal correctness. All experiments use Z3 version 4.15.1, cvc5 version 1.3.1, and Yices version 2.6.4 as the SMT solvers for formal verification.
@@ 4.4 Ablation Study
As demonstrated in Table 3, for problems not solved in the first iteration, the framework’s ability to generate structurally different templates in later iterations proves essential, bringing additional 27 successful barrier synthesis. The second iteration ($I_{2}$) successfully solves 15 additional problems (16.7%), bringing the cumulative success rate to 86.7%. The remaining 12 problems (13.3%) require three or more iterations ($I_{3+}$), with half of these requiring structural refinement ($R_{3-4}$), indicating that more challenging problems often require exploration of alternative mathematical forms, as the proportion of problems requiring structural refinement increases in later iterations (from 7.9% in $I_{1}$ to 50% in $I_{3+}$).
The results of the ablation study provide a positive answer to RQ3, demonstrating the improvement in the efficiency and reliability of LLM-guided barrier synthesis through the agents interaction.
@@ 5 Related Work
Barrier Certificate Generation.
Traditional barrier synthesis has been dominated by optimization-based methods, particularly sum-of-squares (SOS) programming (prajna2007framework; ames2019control; wooding2025protect; kordabad2025sum). To overcome the scalability limitations of SOS, counterexample-guided inductive synthesis (CEGIS) frameworks such as FOSSIL (abate2021fossil) integrate neural function approximators with SMT-based verification, enabling synthesis of nonlinear certificates through iterative counterexample refinement. These approaches substantially expand the range of verifiable systems beyond fixed polynomial templates.
Application of barrier certificates to specifications beyond safety is also studied in the literature (lindemann2018control; jagtap2020formal; majumdar2024necessary). Data-driven variants extend barrier synthesis to systems with unknown or partially known dynamics (nejati2023formal). In the literature on data-driven barrier certificates, most existing methods are restricted to linear or control-affine system dynamics (see, e.g., (Jagtap2020CBCGP; Cohen2022; Lopez2022uCBF)). Moreover, these approaches often depend on known Lipschitz constants to ensure formal guarantees or are applicable only when the system dynamics are partially known. For example, Gaussian processes are used by (Wang2018CBF) and by (Jagtap2020CBCGP) to learn unknown components of nonlinear dynamics while assuming the control-affine part is known. Systems affected by unknown additive disturbances are studied by (Chekan2023UncertainConstraints). Similarly, (mathiesen2024data) assume that the deterministic component of the dynamics is accurately modeled, whereas the barrier-based safety control framework in of (mazouz2024data) presumes known noise distributions. Research addressing systems with fully unknown dynamics are also studied recently (Salamati2021DDCBC; wang2023stochastic; schon2024DRObarrier) with an available software tool LUCID (lucid). Despite these advances, most methods still depend on manual template selection and limited function classes, underscoring the need for adaptive, reasoning-driven synthesis frameworks.
LLMs for Complex Reasoning.
@@
Recent advances in LLMs have demonstrated exceptional reasoning, planning, and problem-solving capabilities (achiam2023gpt; zhao2023survey). Models such as GPT-3 (dale2021gpt3) and GPT-4 (sanderson2023gpt4) exhibit strong generalization and inference abilities by leveraging billions of pretrained parameters and broad world knowledge (xu2024llm). LLMs can decompose complex problems into sub-tasks and reason across multiple abstraction levels, enabling coherent multi-step explanations and decisions (jansen2020visually; lin2023text2motion; rana2023sayplan). These capabilities have been applied across domains—from robotic task planning (rana2023sayplan) and autonomous driving (cui2023large) to mathematical reasoning tasks requiring symbolic manipulation, quantitative analysis, and formal logic (lewkowycz2022solving). Such progress positions LLMs as promising tools for safety-critical domains that demand rigorous mathematical reasoning about system constraints and safe execution.
LLM for CPS.
LLMs are increasingly transforming CPS by introducing high-level reasoning, planning, and natural-language interaction capabilities (xu2024llm). Beyond traditional machine learning models trained on narrow datasets, LLMs draw on diverse web-scale corpora of text and code to serve as both context-aware assistants and autonomous decision-making agents. In robotics, systems such as SayCan (ahn2022saycan) and RT-2 (zitkovich2023rt2) combine LLM reasoning with reinforcement learning or multimodal perception to translate natural-language commands into executable actions. In autonomous driving, DriveGPT4 (yang2023drivegpt4) and Talk2Drive (cui2023large) employ LLMs for interpretable decision-making and natural language interaction.
For formal controller synthesis, bayat2025llm propose a multi-agent framework that translates natural-language specifications into verified symbolic control code, combining reasoning with abstraction-based verification. Despite these advances, integrating LLMs into CPS remains challenging. LLMs are susceptible to hallucination (xu2024llm), lack grounding in physical constraints, and cannot be verified using traditional neural-network methods such as Reluplex due to scale and architectural complexity. These limitations motivate hybrid approaches that couple LLM reasoning with formal verification to ensure correctness and safety in real-world CPS.
6 Limitations
Although our synthesis framework outperforms the baselines by solving 90% of the BarrierBench problems, the remaining 10 unsolved problems reveal some synthesis challenges. The unsolved cases are highly complex nonlinear dynamics, including systems with complex combination of trigonometric and exponential terms, across different topological sets. Their complexity challenges both the synthesis framework and the verification solvers. In certain cases, the Synthesis Agent fails to generate barrier certificates satisfying all safety conditions due to the complexity of the underlying dynamics. Moreover, for the synthesized barriers, SMT solvers face computational difficulties during verifying conditions, frequently timing out or failing to determine satisfiability.

### v2

@@ Abstract
Safety verification of dynamical systems via barrier certificates is central to ensuring correctness in autonomous and other safety-critical applications. Synthesizing such certificates requires discovering mathematical functions that characterize inductive state invariants and provably separate safe and unsafe regions. Existing approaches, however, often struggle to scale computationally, depend on carefully designed templates, and rely on exhaustive or incremental searches over function spaces. They also demand substantial manual ingenuity and mathematical sophistication in constructing the search infrastructure, including selecting template families, choosing appropriate solvers, tuning hyperparameters in data-driven methods, and designing effective sampling procedures.
As a result, successful barrier certificate synthesis requires both a deep understanding of dynamical systems and control theory and practical experience with existing synthesis techniques. Much of this expertise has traditionally been transmitted among practitioners through natural language rather than formalized mathematical procedures.
This observation raises a natural question: can the linguistic and analogical reasoning that experts use informally be captured and operationalized by large language models (LLMs)? Motivated by this question, we present an LLM-agentic framework for barrier certificate synthesis that uses natural language reasoning to propose, refine, and validate candidate certificates. The framework combines LLM-driven template discovery with SMT-based verification and supports barrier-controller co-synthesis for controlled systems, thereby ensuring mathematical compatibility between safety certificates and feedback control laws.
To evaluate this capability, we introduce BarrierBench, a benchmark of 100 dynamical systems spanning linear, nonlinear, discrete-time, and continuous-time settings, including 68 controlled systems that require barrier-controller co-synthesis. Our experiments assess not only the effectiveness of LLM-guided barrier synthesis but also the value of retrieval-augmented generation (RAG) and agentic coordination strategies in improving reliability and performance. Across these tasks, the framework achieves over 90% success in generating valid certificates and demonstrates structural diversity, ranging from simple quadratic forms to high-order coupled polynomials. By releasing BarrierBench with the accompanying toolchain, we aim to establish a community testbed for advancing the integration of language-based reasoning with formal verification for dynamical systems. The benchmark is publicly available at https://hycodev.com/dataset/barrierbench.
@@ 1 Introduction
This century has witnessed an increasingly sophisticated integration of digital software with physical infrastructure, interconnected through wired and wireless networks. These cyber-physical systems (CPS) operate so seamlessly and intuitively that it is easy to overlook the complexity beneath their surface: their successful operation depends on intricate control logic whose correctness requires a deep understanding of both the underlying physics and the discrete switching logic governing their behavior. As technological practice continues to deliver critical infrastructures such as autonomous vehicles, implantable medical devices, and smart grids, ensuring their safety has become an increasingly demanding challenge for control engineers amid ever-growing system complexity. In this paper, we address this challenge by exploring how large language models (LLMs) can assist in the synthesis and verification of safety guarantees for complex CPS.
@@
Safety verification of dynamical systems is a canonical problem in the design of safety-critical CPS. Since safety verification is computationally undecidable even for CPS with simple piecewise-constant dynamics (Asarin and Maler, 1998), fully automatic procedures for obtaining provable safety guarantees remain out of reach. Consequently, a variety of methods has been developed to construct safety proofs, including symbolic state-space exploration, statistical guarantees, and abstraction-based techniques. Among these, the synthesis of barrier certificates (Prajna and Jadbabaie, 2004; Ames et al., 2019; Jagtap et al., 2021) has emerged as a particularly effective and scalable approach. Barrier certificates serve as functional analogs of inductive state invariants, characterizing the separation between safe and unsafe regions of the state space. Originating in Lyapunov’s classical theory of stability (Lyapunov, 1966; Khalil and Grizzle, 2002), they extend the central idea of certifying system behavior from convergence to invariance. Their formulation as the search for a single scalar function satisfying suitable inequality constraints makes them amenable to optimization- and satisfiability-based methods, contributing to their prominence in both theory and practice. We therefore focus on barrier-certificate-based proof discovery as a promising foundation for integrating formal reasoning with language-model-guided synthesis.
@@
A barrier certificate is a scalar function whose level sets define an inductive invariant separating unsafe states from all trajectories originating from admissible initial conditions. By providing formal guarantees over infinite time horizons, barrier certificates are central to correctness arguments for both linear and nonlinear dynamical systems. Traditional approaches to barrier synthesis are predominantly optimization-based. Sum-of-squares and semidefinite programming methods cast the search as a convex optimization problem over polynomial templates (Prajna et al., 2007; Kordabad et al., 2025). These methods provide strong correctness guarantees, but their computational cost grows rapidly with system dimension, and they are typically limited to low-degree polynomial classes. To improve generality, counterexample-guided inductive synthesis (CEGIS) frameworks—see Dawson et al. (2023) for a comprehensive review—combine such methods with Satisfiability Modulo Theories (SMT) solvers to iteratively refine candidate barriers (Ravanbakhsh and Sankaranarayanan, 2015). Neural and data-driven variants use deep architectures to approximate more expressive certificate functions (Abate et al., 2024; Neustroev et al., 2025; Abate et al., 2021). Although these approaches broaden the space of representable barriers, they also introduce new challenges, including training instability, reliance on finite samples, reduced interpretability, and continued dependence on carefully chosen templates or architectures. As a result, successful synthesis still depends heavily on expert intuition in selecting suitable function classes, solvers, and hyperparameters. These decisions often reflect informal reasoning that remains difficult to encode within existing algorithmic frameworks. This paper investigates whether LLMs can capture and operationalize the linguistic and analogical reasoning that practitioners use when constructing barrier certificates.
@@
We present an LLM-guided synthesis framework for barrier certificates that leverages natural-language reasoning to propose, refine, and validate candidate certificates. The framework combines LLM-driven template discovery with SMT-based formal verification and naturally extends to barrier–controller co-synthesis for controlled systems. By incorporating verification feedback into subsequent LLM prompts, the framework supports dynamic exploration of barrier templates and coefficients, often uncovering novel mathematical forms that fixed-template approaches (Prajna et al., 2007; Kong et al., 2013; Ames et al., 2019; Abate et al., 2021; Kordabad et al., 2025) may overlook. LLMs guide the synthesis process by reasoning about system properties, safety requirements, and mathematical structure, first analyzing local safety conditions and then integrating them into global certificates. This integration provides the expressiveness and adaptability needed for complex systems while preserving the mathematical rigor required in safety-critical applications.
Figure 1: Agentic framework for LLM-guided barrier certificate synthesis.
@@
To operationalize this idea, we design an agentic architecture that emulates the workflow of expert control designers. Human designers typically draw on prior experience with similar dynamical systems to select barrier templates, solver strategies, and proof heuristics. Analogously, our framework instantiates three collaborating agents: a Barrier Retrieval Agent, which searches a database of previously solved systems to identify analogous examples; a Barrier Synthesis Agent, which proposes candidate barrier certificates inspired by retrieved analogs; and a Barrier Verifier Agent, which evaluates these candidates in two stages, first through a lightweight data-driven check and then through formal SMT-based verification. Feedback from the Verifier Agent is returned to the Synthesis Agent, forming a closed-loop interaction bounded by a global timeout. This design enables iterative refinement of candidate certificates while making effective use of external formal tools. Our study investigates the following research questions:
@@ 3. RQ3:
To evaluate these questions, we introduce BarrierBench, a benchmark of 100 dynamical systems spanning linear, nonlinear, discrete-time, and continuous-time settings, including 68 controlled systems that require barrier–controller co-synthesis. Across these tasks, the LLM framework achieves over 90% success in generating valid certificates and exhibits structural diversity ranging from simple quadratics to high-order coupled polynomials. By releasing BarrierBench and the accompanying toolchain, we aim to establish a community testbed for advancing the integration of language-based reasoning with formal verification for dynamical systems.
@@ 2 Barrier Certificates
Theorem 1 (Safety via Barrier Certificates (Prajna and Jadbabaie, 2004)).
@@
The above definitions and theorem are stated for systems without control inputs. If the system has control input $u\in U$, the objective is to design a state feedback law that gives $u$ as a function of $x$ and makes the system safe. For this, the barrier conditions are quantified existentially over the input.
@@
 $\text{e}\leftarrow\textsc{RetrievalAgent}(f,X_{I},X_{U},\mathcal{D})$; $(B_{k},C_{k})\leftarrow\textsc{SynthesisAgent}(\text{e},\textsf{feedback}_{k})$
@@ 3 Agentic Framework for Barrier Synthesis
Our synthesis framework employs a multi-agent architecture in which specialized LLM-powered agents collaborate to discover barrier certificates. The framework comprises three core agents operating in an iterative pipeline: (1) a Barrier Retrieval Agent that identifies relevant prior solutions; (2) a Barrier Synthesis Agent that generates candidate barrier certificates; and (3) a Barrier Verifier Agent that checks candidates and provides feedback for refinement. This iterative process addresses limitations of conventional approaches by enabling informed template selection and adaptive exploration of diverse mathematical structures through verification-guided synthesis.
@@
The Barrier Verifier Agent conducts a two-stage evaluation of candidate barriers. In the first stage, it performs a sample-based validation across $N$ points drawn from $X_{I}$, $X_{U}$, and $X$, where $N$ is selected to balance computational efficiency with adequate state space coverage; we use $N=5000$ in our experiments to achieve this trade-off, discarding invalid barriers that violate any of the three conditions on at least one of sampled points, to avoid unnecessary SMT solver calls. In the second stage, it performs formal verification using SMT solvers such as Z3 (de Moura and Bjørner, 2008), Yices (Dutertre, 2014), and cvc5 (Barbosa et al., 2022), checking the unsatisfiability of negated barrier conditions. The agent selects solvers adaptively based on problem structure (e.g., polynomial vs. transcendental terms), expression complexity, and implicit knowledge of solver capabilities (e.g., whether transcendental functions are supported), and then applies symbolic conversions or Taylor approximations when necessary, and handles verification failures through solver switching or adaptive timeout adjustments.
@@ 4.1 Experimental Setup
The framework operates with carefully chosen hyperparameters validated across various systems. We set maximum main iterations $K=5$ and maximum refinements per iteration $R=4$. The LLM utilizes Claude Sonnet 4 and ChatGPT-4o with engineered prompts designed to synthesize mathematical expressions while maintaining formal correctness. All experiments use Z3 version 4.15.1, cvc5 version 1.3.1, and Yices version 2.6.4 as the SMT solvers for formal verification. For the verifier agent, we use in total $N=5000$ samples with 30% from the initial set, 30% from the unsafe set, and 40% from the state space.
@@ 4.4 Ablation Study
Our retrieval does not transfer certificates directly. It provides structural templates as few-shot examples for the LLM. Similar to few-shot learning (Brown et al., 2020), the synthesis agent uses architectural patterns from analogous problems, then generates entirely new coefficients and controllers tailored to the current problem’s specific constraints while explicitly instructed not to copy solutions (see Appendix A). For out-of-distribution scenarios where retrieval fails, the framework gracefully degrades to independent synthesis (baseline 41% capability) with iterative structural exploration. Importantly, as discussed in Section 3, failures from initial attempts accumulate as feedback, constrain the exploration space by showing what does not work, enabling the synthesis agent to propose progressively better structures in subsequent iterations.
@@
As demonstrated in Table 3, for problems not solved in the first iteration, the framework’s ability to generate structurally different templates in later iterations proves essential, bringing additional 27 successful barrier synthesis. The second iteration ($I_{2}$) successfully solves 15 additional problems (16.7%), bringing the cumulative success rate to 86.7%. The remaining 12 problems (13.3%) require three or more iterations ($I_{3+}$), with half of these requiring structural refinement ($R_{3-4}$), indicating that more challenging problems often require exploration of alternative mathematical forms, as the proportion of problems requiring structural refinement increases in later iterations (from 7.9% in $I_{1}$ to 50% in $I_{3+}$). The results of the ablation study provide a positive answer to RQ3, demonstrating the improvement in the efficiency and reliability of LLM-guided barrier synthesis through the agents interaction.
Limitations.
Although our synthesis framework outperforms the baselines by solving 90% of the BarrierBench problems, the remaining 10 unsolved problems reveal some synthesis challenges. The unsolved cases are highly complex nonlinear dynamics, including systems with complex combination of trigonometric and exponential terms, across different topological sets. Their complexity challenges both the synthesis framework and the verification solvers. In certain cases, the Synthesis Agent fails to generate barrier certificates satisfying all safety conditions due to the complexity of the underlying dynamics. Moreover, for the synthesized barriers, SMT solvers face computational difficulties during verifying conditions as system dimension and nonlinear complexity increase, frequently timing out or failing to determine satisfiability.
@@ 5 Related Work
Barrier certificate generation. Classical approaches to barrier certificate synthesis are predominantly optimization-based, especially sum-of-squares (SOS) programming (Prajna et al., 2007; Ames et al., 2019; Wooding et al., 2026; Kordabad et al., 2025), but their computational cost grows quickly with system dimension and complexity. To improve scalability, counterexample-guided inductive synthesis (CEGIS) frameworks such as FOSSIL (Abate et al., 2021) integrate neural networks with SMT-based verification, enabling iterative refinement of nonlinear barrier certificates beyond fixed polynomial templates. Barrier certificates have also been extended beyond safety to richer specifications, including signal temporal logic (Lindemann and Dimarogonas, 2019), linear temporal logic over finite traces (Jagtap et al., 2021), and reachability specifications (Majumdar et al., 2024).
When system dynamics are difficult to model exactly, data-driven approaches aim to infer safe behavior from sampled trajectories (Nejati et al., 2023). Much of this literature focuses on linear or control-affine dynamics (Jagtap et al., 2020; Cohen et al., 2022; Lopez and Slotine, 2023), often under partial knowledge or structural assumptions needed for formal guarantees. For example, (Wang et al., 2018) learns unknown dynamics with Gaussian processes while assuming known control-affine structure, and (Jagtap et al., 2020) assumes known control input while learning the unknown drift. Other works impose assumptions on stochastic components, for example by assuming unknown noise distributions over known deterministic dynamics (Mathiesen et al., 2024) or known noise distributions (Mazouz et al., 2024). More recent work considers systems with fully unknown dynamics (Salamati et al., 2024; Wang et al., 2023; Schön et al., 2024), supported by tools such as LUCID (Casablanca et al., 2026). Despite these advances, most existing methods still rely on manually chosen templates, motivating more flexible and automated synthesis approaches.
@@
LLMs for complex reasoning. Recent LLMs exhibit strong reasoning, planning, and problem-solving abilities (Achiam et al., 2023; Xu et al., 2024a; Zhao et al., 2023). Models such as GPT-3 (Brown et al., 2020; Dale, 2021) and GPT-4 (Sanderson, 2023) demonstrate broad generalization grounded in large-scale pretraining (Xu et al., 2024a). They can generate multi-step plans from high-level instructions (Jansen, 2020), decompose complex problems into subproblems, and reason across multiple levels of abstraction (Lin et al., 2023; Rana et al., 2023). These capabilities have been applied in domains ranging from robotic task planning (Rana et al., 2023) and autonomous driving (Cui et al., 2024) to mathematical reasoning involving symbolic manipulation, quantitative analysis, and formal logic (Lewkowycz et al., 2022). This progress makes LLMs promising tools for safety-critical settings that require structured reasoning about system constraints and safe behavior.
LLMs for CPS. LLMs are increasingly being explored in CPS as engines for high-level reasoning, planning, and natural-language interaction (Xu et al., 2024a). Their broad pretraining across language, code, mathematics, and science gives them a wide base of transferable knowledge. In robotics, SayCan (Ichter et al., 2023) combines LLM reasoning with reinforcement learning by using affordance-based value functions to ground language commands in executable actions. In autonomous driving, DriveGPT4 (Xu et al., 2024b) uses LLMs for interpretable decision-making, generating natural-language explanations together with low-level control predictions. In formal controller synthesis, (Bayat et al., 2025) propose a multi-agent framework that translates natural-language specifications into verified symbolic control code by coupling reasoning with abstraction-based verification. At the same time, LLMs remain prone to hallucination (Xu et al., 2024a), lack intrinsic grounding in physical constraints, and are difficult to verify with traditional neural-network verification tools such as Reluplex. These limitations motivate hybrid frameworks that combine LLM-based reasoning with formal verification to ensure correctness and safety in CPS.